# Gapfree graphs and powers of edge ideals with linear quotients

### v1

@@ 2. Duplication of finite graphs
Let $H=G^{*}$ be the graph obtained from $G$ by duplicating the vertex $x\in V(G)$. Let $y\in V(H)$ be the duplicate of $x$ in $H$. Let $\mathcal{G}(I)$ denote the minimal generating set of $I$. Let $u\in\mathcal{G}(I(G)^{s})$ with $x$-degree $r>0$. Then $u$ will produce $r$ new generators in $\mathcal{G}(I(H)^{s})$, namely
(1) $\frac{\,u\,}{x}y,\frac{\,u\,}{x^{2}}y^{2},\dots,\frac{\,u\,}{x^{r}}y^{r}$
In fact, all of the new generators will be of this form since $x$ and $y$ are non-adjacent in $H$. Extend the linear quotients order on $\mathcal{G}(I(G)^{s})$ to a linear quotients order on $\mathcal{G}(I(H)^{s})$ as follows: If $u\in\mathcal{G}(I(G)^{s})$ has positive $x$-degree $r$, then insert the new generators as in (1) right after $u$.
Let $w_{1},\dots,w_{q}$ be the order on the minimal generators of $I(H)^{s}$. Let $t\geq 2$. We will show that $I=(w_{1},\dots,w_{t-1}):w_{t}$ is generated by variables. Let $\ell<t$ be fixed. We will show that there is a variable generator of $I$ that divides $w_{\ell}:w_{t}$.
(CASE 1) Suppose that $y$ does not divide $w_{t}$. Let $J$ be the ideal generated by $\{w_{i}:w_{t}\mid 1\leq i<t,w_{i}\in\mathcal{G}(I(G)^{s})\}$. Then $J$ is generated by variables and $\mathcal{G}(J)\subseteq\mathcal{G}(I)$. Therefore, we may assume that $y$ divides $w_{\ell}$.
We now claim that $x\in\mathcal{G}(J)$ implies $y\in\mathcal{G}(I)$. Indeed, if $x\in\mathcal{G}(J)$, then $w_{t^{\prime}}:w_{t}=x$ for some $w_{t^{\prime}}\in\mathcal{G}(I(G)^{s})$ with $t^{\prime}<t$. Since $w_{t^{\prime}}$ has positive $x$-degree, by definition of the order, $w_{t^{\prime\prime}}=w_{t^{\prime}}y/x$ for some $t^{\prime}<t^{\prime\prime}<t$. Then $w_{t^{\prime\prime}}:w_{t}=y$ and $y\in\mathcal{G}(I)$ as wanted.
Case 1.1: If $x\in\mathcal{G}(J)$, then by the above claim $y\in\mathcal{G}(I)$, and thus $y$ divides $w_{\ell}:w_{t}$.
Case 1.2: Suppose that $x\notin\mathcal{G}(J)$. Let $w_{\ell}=x^{q_{1}}y^{q_{2}}v$ where $q_{1}\geq 0$ is the $x$-degree of $w_{\ell}$ and $q_{2}\geq 1$ is the $y$-degree of $w_{\ell}$. By definition of the order, $w_{\ell^{\prime}}=x^{q_{1}+q_{2}}v$ for some $\ell^{\prime}<\ell$. Then, there is a variable generator $z\neq x$ of $J$ such that $z$ divides $w_{\ell^{\prime}}:w_{t}$. This implies that $z$ divides $v:w_{t}$. Thus, $z$ divides $w_{\ell}:w_{t}$ as desired.
(CASE 2) Suppose that $y$ divides $w_{t}$. First, we claim that $x\in\mathcal{G}(I)$. Indeed, by definition of the order, $w_{\eta}=xw_{t}/y$ for some $\eta<t$ and thus $w_{\eta}:w_{t}=x\in\mathcal{G}(I)$.
Let $w_{\ell}=x^{\gamma}y^{\sigma}N$ where $\gamma$ is the $x$-degree, and $\sigma$ is the $y$-degree of $w_{\ell}$. Let $w_{t}=x^{\alpha}y^{\beta}M$ where $\alpha$ is the $x$-degree, and $\beta\geq 1$ is the $y$-degree of $w_{t}$. Let $w_{L}=x^{\gamma+\sigma}N$ and $w_{T}=x^{\alpha+\beta}M$. Then, by definition of the order, $L<\ell<T<t$.
If $\gamma>\alpha$, then nothing is left to show as $x\in\mathcal{G}(I)$. Therefore, let us assume that $\gamma\leq\alpha$.
Case 2.1: Suppose that $\gamma+\sigma\leq\alpha+\beta$. Since $I(G)^{s}$ has linear quotients, there exists $w_{T^{\prime}}\in\mathcal{G}(I(G)^{s})$ with $T^{\prime}<T$ such that $w_{T^{\prime}}:w_{T}=z$ for some variable $z$ that divides $w_{L}:w_{T}=N:M$. Then $z$ divides $w_{\ell}:w_{t}$. Note that $z\neq x,y$. It remains to show that $z$ is a generator of $I$. Since $w_{T}$ and $w_{T^{\prime}}$ are of the same degree, and $w_{T^{\prime}}:w_{T}=z$, either
  • (a)
    $w_{T^{\prime}}=x^{\alpha+\beta-1}Mz$ or
  • (b)
    $w_{T^{\prime}}=x^{\alpha+\beta}Mz/z^{\prime}$ for some variable $z^{\prime}(\neq z)$ dividing $M$.
If (a) holds, then $w_{t^{\prime}}=x^{\alpha}y^{\beta-1}Mz$ for some $t^{\prime}$ that satisfies $T^{\prime}\leq t^{\prime}<T<t$. Since $w_{t^{\prime}}:w_{t}=z$, it follows that $z\in\mathcal{G}(I)$ as wanted.
If (b) holds, then $w_{t^{\prime}}=x^{\alpha}y^{\beta}Mz/z^{\prime}$ for some $t^{\prime}$ that satisfies $T^{\prime}<t^{\prime}<T<t$. Since $w_{t^{\prime}}:w_{t}=z$, it follows that $z\in\mathcal{G}(I)$ as wanted.
Case 2.2: Suppose that $\gamma+\sigma>\alpha+\beta$. Since we are also assuming that $\gamma\leq\alpha$, it follows that $\sigma>\beta$. Then $w_{\ell}:w_{t}=y^{\sigma-\beta}(N:M)$. Then, it remains to show that either $y\in\mathcal{G}(I)$, or there is a variable in $\mathcal{G}(I)$ that divides $N:M$.
Since $I(G)^{s}$ has linear quotients, there exists $w_{L^{\prime}}\in\mathcal{G}(I(G)^{s})$ with $L^{\prime}<T$ such that $w_{L^{\prime}}:w_{T}=z$ for some variable $z$ that divides $w_{L}:w_{T}=x^{\gamma+\sigma-(\alpha+\beta)}(N:M)$.
Case 2.2.1: Suppose that $z=x$. Then, $w_{L^{\prime}}=x^{\alpha+\beta+1}M/z^{\prime}$ for some variable $z^{\prime}$ dividing $M$. Then $w_{\ell^{\prime}}=x^{\alpha}y^{\beta+1}M/z^{\prime}$ for some $\ell^{\prime}$ that satisfies $L^{\prime}<\ell^{\prime}<T<t$. It follows that $w_{\ell^{\prime}}:w_{t}=y$ and $y\in\mathcal{G}(I)$ as desired.
Case 2.2.2: Suppose that $z\neq x$, and $z$ divides $N:M$. It suffices to show that $z\in\mathcal{G}(I)$. Since $w_{L^{\prime}}$ and $w_{T}$ have the same degree, either
  • (a)
    $w_{L^{\prime}}=x^{\alpha+\beta-1}Mz$ or
  • (b)
    $w_{L^{\prime}}=x^{\alpha+\beta}Mz/z^{\prime}$ for some variable $z^{\prime}(\neq z)$ dividing $M$.
If $(a)$ holds, then $w_{\ell^{\prime}}=x^{\alpha}y^{\beta-1}Mz$ for some $\ell^{\prime}$ with $L^{\prime}\leq\ell^{\prime}<T<t$. Then, $w_{\ell^{\prime}}:w_{t}=z$ and $z\in\mathcal{G}(I)$ as desired.
If $(b)$ holds, then $w_{\ell^{\prime}}=x^{\alpha}y^{\beta}Mz/z^{\prime}$ for some $\ell^{\prime}$ with $L^{\prime}<\ell^{\prime}<T<t$. Then, $w_{\ell^{\prime}}:w_{t}=z$ and $z\in\mathcal{G}(I)$ as desired.
@@ 3. Ordering of pure powers
Let $G$ be a gapfree graph with $E(G)=\{e_{1},\ldots,e_{s}\}$ and suppose that a power $I(G)^{q}$ of the edge ideal $I(G)$ has linear quotients with an ordering
${\mathcal{N}}:u_{1}>\cdots>u_{r}$
of the generators of $I(G)^{q}$, where pure powers appear to be
@@
${\mathcal{N}}^{(1)}:u_{1}e_{1}>\cdots>u_{r}e_{1}>u_{1}e_{2}>\cdots>u_{r}e_{2}>\cdots>u_{r}e_{s}$
of the generators of $I(G)^{q+1}$, where $u_{i}e_{j}$ is omitted if $u_{i}e_{j}=u_{k}e_{\ell}$ with $\ell<j$, is called the efficient ordering coming from the ordering ${\mathcal{N}}$. Recurrently we define the ordering ${\mathcal{N}}^{(j)}$ of the generators of $I(G)^{q+j}$ for $j\geq 2$ by setting
${\mathcal{N}}^{(j)}=({\mathcal{N}}^{(j-1)})^{(1)}.$
Strong evidence for Conjecture 1.2 is the following:
@@
Let $G$ be a gapfree graph with $E(G)=\{e_{1},\ldots,e_{s}\}$ for which a power of $I(G)$ has linear quotients. Let $q_{0}$ denote the smallest integer for which $I(G)^{q_{0}}$ has linear quotients. Suppose that $I(G)^{q_{0}}$ has linear quotients with an ordering ${\mathcal{N}}:u_{1}>\cdots>u_{r}$ of the generators of $I(G)^{q_{0}}$. Then each power $I(G)^{q}$ for $q>q_{0}$ has linear quotients with the ordering ${\mathcal{N}}^{(q-q_{0})}$ of the generators of $I(G)^{q}$.
@@
Let ${\mathcal{A}}(e_{1})$ denote the set of those $w_{i}$ which is divided by $e_{1}$, ${\mathcal{A}}(e_{2})$ the set of those $w_{i}\not\in{\mathcal{A}}(e_{1})$ which is divided by $e_{2}$, ${\mathcal{A}}(e_{3})$ the set of those $w_{i}\not\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$ which is divided by $e_{3}$ and ${\mathcal{A}}(e_{5})$ the set of those $w_{i}\not\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$ which is divided by $e_{5}$. Finally, ${\mathcal{A}}(e_{4})=\{e_{4}^{q}\}$.
We claim that the colon ideal $(w_{1},\ldots,w_{p-1}):w_{p}$ is generated by a subset of $\{a,b,c,d,e\}$. If $w_{p}\in{\mathcal{A}}(e_{1})$, then the claim follows from the induction hypothesis.
Let $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{2})$ with $i<p$. Let $w_{i}=u_{s}e_{2}$ and $w_{p}=u_{t}e_{2}$ with $s<t$. Then there is $u_{\ell}$ with $\ell<t$ for which $u_{\ell}:u_{t}$ is a variable and $(u_{\ell}:u_{t})|(u_{s}:u_{t})$. One has $j<p$ with $w_{j}=u_{\ell}e_{2}$. Note that $w_{j}$ may belong to ${\mathcal{A}}(e_{1})$. Then $w_{j}:w_{p}$ is a variable and $(w_{j}:w_{p})|(w_{i}:w_{p})$, as desired.
Let $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{1})$. If $a$ does not appear in $w_{p}$, then $w_{i}:w_{p}$ is divided by $a=(e_{1}w_{p}/e_{2}):w_{p}$. If $e_{5}=ae$ divides $w_{p}$, then both $(e_{1}w_{p}/e_{5}):w_{p}=b$ and $(e_{1}e_{3}w_{p}/e_{2}e_{5}):w_{p}=d$ belong to $(w_{1},\ldots,w_{p-1}):w_{p}$. Let $w_{i}=e_{1}^{f_{1}}e_{2}^{f_{2}}e_{3}^{f_{3}}e_{4}^{f_{4}}e_{5}^{f_{5}}$ and $w_{p}=e_{2}^{f^{\prime}_{2}}e_{3}^{f^{\prime}_{3}}e_{4}^{f^{\prime}_{4}}e_{5}^{f^{\prime}_{5}}$. If one of the inequalities
@@
which is impossible, since $f_{1}>0$.
Let $w_{p}\in{\mathcal{A}}(e_{3})$. Since $e_{2}(w_{p}/e_{3})\in{\mathcal{A}}(e_{2})$, it follows that $(e_{2}w_{p}/e_{3}):w_{p}=b$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $b$ does not appear in $e_{p}$, if $e_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$, then $b$ divides $w_{i}:w_{p}$. If $e_{i}\in{\mathcal{A}}(e_{3})$, then the similar argument as in the case of $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{2})$ is valid without modification.
Let $w_{p}\in{\mathcal{A}}(e_{5})$. Since $e_{1}(w_{p}/e_{5})\in{\mathcal{A}}(e_{1})$, it follows that $(e_{1}w_{p}/e_{5}):w_{p}=b$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $b$ does not appear in $e_{p}$, if $e_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$, then $b$ divides $w_{i}:w_{p}$. If $e_{4}$ divides $w_{p}$, then $e_{3}(w_{p}/e_{4})\in{\mathcal{A}}(e_{3})$ and $(e_{3}w_{p}/e_{4}):w_{p}=c$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $c$ does not appear in $e_{p}$, if $e_{i}\in{\mathcal{A}}(e_{3})$, then $c$ divides $w_{i}:w_{p}$. Let $w_{p}=e_{5}^{q}$. Since $e_{4}e_{5}^{q-1}>e_{5}^{q}$ and $e_{4}e_{5}^{q-1}:e_{5}^{q}=d$ belongs $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $d$ does not appear in $w_{p}=e_{5}^{q}$, if $e_{i}\in{\mathcal{A}}(e_{3})$, then $d$ divides $w_{i}:w_{p}$. If $e_{i}\in{\mathcal{A}}(e_{5})$, then the similar argument as in the case of $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{2})$ is valid without modification.
Finally, if $w_{p}\in{\mathcal{A}}(e_{4})$, in other words, $e_{p}=e_{4}^{q}$, then a routine computation done as above yields $(w_{1},\ldots,w_{p-1}):w_{p}=(a,c)$. This completes the proof.
@@
4. Pentagon together with vertices
Developing the argument done in the proof of Proposition 3.2, we now study finite graphs formed by carefully adding vertices to the pentagon. We first study the case where a single vertex is added with specific connecting edges.

### v2

@@ 3. Expansion, and linear quotients of powers of edge ideals
First note that if $I(G^{[x]})^{s}$ has linear quotients for some $s\geq 1$ then $G^{[x]}$ is gapfree, thus we focus on the converse.
For any monomial generator $w$ of $I(G^{[x]})^{s}$ define
$\mu(w)=\min\left\{i:\frac{w}{(xy)^{i}}\in\mathcal{G}(I(G^{x})^{s-i})\right\}.$
Note that $\mu(w)=0$ if and only if $w$ is a minimal generator of $I(G^{x})^{s}$. Set $V=V(G^{[x]})$. Partition $V$ into sets:
$Z$ $=$ $\{x,y\}$
$A$ $=$ $N_{G}(x)$
$B$ $=$ $V(G^{[x]})\setminus N_{G^{[x]}}[x]$
By 3.3, $B$ is an independent set. Fix an order $>_{B}$ on the elements of $B$. By Proposition 2.3, there is a linear quotients order $w_{1},w_{2},\dots,w_{q}$ for $I(G^{x})^{s}$. We extend this order to a linear quotients order $w_{1},\dots,w_{q},w_{q+1},\dots,w_{r}$ for $I(G^{[x]})^{s}$. To simplify the notation, we adopt the convention that $w_{1}>w_{2}>\dots>w_{r}$. For generators $w$ and $w^{\prime}$ of $I(G^{[x]})^{s}$ not belonging to $I(G^{x})^{s}$, set $w>w^{\prime}$ if:
  1. (1)
    $\mu(w)<\mu(w^{\prime})$, or
  2. (2)
    $\mu(w)=\mu(w^{\prime})$ and $\deg(w_{Z})<\deg(w^{\prime}_{Z})$, or
  3. (3)
  4. (4)
    $\mu(w)=\mu(w^{\prime})$, $\deg(w_{Z})=\deg(w^{\prime}_{Z})$, $|\deg_{x}(w)-\deg_{y}(w)|=|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|$, and $w_{B}>w^{\prime}_{B}$ in the lexicographic ordering on $B$, using the order $>_{B}$ on the elements of $B$ fixed above.
Let $w=w_{i}$ for some $i\geq 2$. We show that $J=(w_{1},\dots,w_{i-1}):w_{i}$ is generated by variables. We may assume that $\mu(w)\neq 0$ by the definition of the order and Proposition 2.3. Since $B$ is an independent set, the edges of $G^{[x]}$ can be partitioned as follows:
$E_{1}$ $=$ $\{xy\}$
$E_{2}$ $=$ $\{ab\mid a\in A,b\in B\}$
$E_{3}$ $=$ $\{az\mid a\in A,z\in Z\}$
$E_{4}$ $=$ $\{aa^{\prime}\mid a,a^{\prime}\in A\}.$
We make two observations which will be used in the proofs.
  • (i)
    Let $w$ be a generator of $I(G^{[x]})^{s}$ with $\mu(w)\neq 0$. For any representation
    $w=(xy)^{\mu(w)}e_{1}\dots e_{s-\mu(w)}$
    where $e_{1},\dots,e_{s-\mu(w)}$ are in $E(G^{x})$, we have $e_{j}\notin E_{4}$ for all $j=1,\dots,s-\mu(w)$.
  • (ii)
    $A$ is a subset of the set of generators of $J$.
$w=(xy)^{\mu(w)-1}(xa_{1})(ya_{2})e_{2}\dots e_{s-\mu(w)}$
which is a contradiction to the minimality of $\mu(w)$.
Proof of (ii). Let $a\in A$. Then $u:=\frac{w}{xy}(xa)$ is a minimal generator of $I(G^{[x]})^{s}$ with $\mu(u)\leq\mu(w)-1$. Therefore $u>w$ and $u:w=a\in J$, proving the claim.
Let $v>w$. We show that there is a variable generator of $J$ dividing $v:w$.
$\deg(v_{A})\geq s-\mu(v)>s-\mu(w)=\deg(w_{A}).$
$\deg(v_{A})=s-\mu(v)=s-\mu(w)=\deg(w_{A}).$
$2\gamma+m=\deg(v_{Z})<\deg(w_{Z})=2\gamma+m^{\prime}.$
Then since $v_{A}=w_{A}$, there exists $a\mid(f_{1}\dots f_{k})_{A}$ such that $a\mid(g_{1}\dots g_{m^{\prime}})_{A}$. Without loss of generality, we may assume that $f_{1}=ab$ for some $b\in B$ and $g_{1}=az$ for some $z\in Z$. If $b\mid(v:w)$, set $w^{\prime}:=\frac{w}{g_{1}}f_{1}=\frac{w}{z}b$. Then $w^{\prime}>w$ with $w^{\prime}:w=b$, which completes the proof.
Case 2.2: Suppose that $\deg(v_{Z})=\deg(w_{Z})$.
Case 2.2-(i): Suppose that $|\deg_{x}(v)-\deg_{y}(v)|<|\deg_{x}(w)-\deg_{y}(w)|$. Since the right side of the inequality cannot be 00, without loss of generality, assume that $\deg_{x}(w)>\deg_{y}(w)$. Then since $\deg(v_{Z})=\deg(w_{Z})$, we have $\deg_{y}(v)>\deg_{y}(w)$ and $y$ divides $v:w$.
Since $\deg_{x}(w)>\deg_{y}(w)$ there is an $a\in A$ with $e_{1}=ax$ in the representation of $w$. By the definition of $G^{[x]}$, $a\in N_{G^{[x]}}(y)$ also. Set $w^{\prime}:=\frac{w}{ax}ay$. If $|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|=|\deg_{x}(w)-\deg_{y}(w)|$, then $|\deg_{x}(w)-\deg_{y}(w)|=1$ and $|\deg_{x}(v)-\deg_{y}(v)|=0$. In this case, $\deg(w_{Z})$ is odd and $\deg(v_{Z})$ is even, which contradicts $\deg(v_{Z})=\deg(w_{Z})$. Thus $|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|<|\deg_{x}(w)-\deg_{y}(w)|$, so $w^{\prime}>w$ and $w^{\prime}:w=y$.
Case 2.2-(ii): Suppose that $|\deg_{x}(v)-\deg_{y}(v)|=|\deg_{x}(w)-\deg_{y}(w)|$. Assume first that $v_{Z}\neq w_{Z}$. Then without loss of generality, we may write
  • $\deg_{x}(v)=r_{1}=\deg_{y}(w)$
  • $\deg_{x}(w)=r_{2}=\deg_{y}(v)$
for some integers $r_{1}>r_{2}$. Then there exists an $e=ay\in E_{3}$ in the representation of $w$. Set $w^{\prime}:=\frac{w}{ay}ax=\frac{w}{y}x$. Then $w^{\prime}>w$ and $x=w^{\prime}:w$ divides $v:w$. Therefore, we may assume $v_{Z}=w_{Z}$. Recall that the following statements are now assumed:
  • $1\leq\mu(v)=\mu(w)=\gamma$,
  • $v_{A}=w_{A}$,
  • $v_{Z}=w_{Z}$.
Fix presentations $v=(xy)^{\gamma}f_{1}f_{2}\cdots f_{t}$ and $w=(xy)^{\gamma}g_{1}g_{2}\cdots g_{t}$. For simplicity, we assume that $f_{i_{1}}\dots f_{i_{m}}\neq g_{j_{1}}\dots g_{j_{m}}$ for all $m$-fold products with $1\leq m\leq t$ since the argument below remains valid recursively after removing the common factors.
Fix $p$ be the greatest in the ordering fixed on $B$ such that $p\mid(v:w).$ We claim that there is $u>w$ such that $u:w=p$. After rearranging the indices, we may write $f_{1}=a_{1}p$ for some $a_{1}\in A$.
Step 1: If there exists $\delta_{1}\in Z$ or $\delta_{1}\in B$ with $p>_{B}\delta_{1}$ such that $g_{1}=a_{1}\delta_{1}$ after rearrangement of indices, set
$u=\frac{w}{g_{1}}f_{1}=\frac{w}{(a_{1}\delta_{1})}(a_{1}p)$
Then $u>w$ in the ordering since $\deg(u_{Z})<\deg(w_{Z})$ in the first case and $u_{B}>w_{B}$ in the latter one, and $u:w=p$. Therefore, we may assume that $g_{1}=a_{1}\delta_{1}$ for some $\delta_{1}\in B$ with $\delta_{1}>_{B}p$.
Step 2: By the choice of $p$, we may set $f_{2}=a_{2}\delta_{1}$ for some $a_{2}\in A$. As before, if, up to reordering, $g_{2}=a_{2}\delta_{2}$ for some $\delta_{2}\in Z$ or $\delta_{2}\in B$ with $p>_{B}\delta_{2}$ we form the monomial
which satisfies the desired property. Therefore, we may assume that $\delta_{2}\in B$ and $\delta_{2}>_{B}p$.
Step 3: By the choice of $p$, we may set $f_{3}=a_{3}\delta_{2}$ for some $a_{3}\in A$. As before, if, up to reordering, $g_{3}=a_{3}\delta_{3}$ for some $\delta_{3}\in Z$ or $\delta_{3}\in B$ with $p>_{B}\delta_{3}$ we form the monomial
which satisfies the desired property. Therefore, we may assume $\delta_{3}\in B$ and $\delta_{3}>_{B}p$. The process must stop after $k\leq t$ steps when there is no $\delta_{k}\in B$ left with $\delta_{k}>_{B}p$. When the process stops, we have $u>w$ and $u:w=p$ as desired. ∎
Corollary 3.5.
Let $G$ be a graph and assume $I(G)^{s}$ has linear quotients for some $s$. Let $x$ be a vertex of $G$ such that $V(G)\setminus N_{G}[x]$ is an independent set. Let $H$ be the graph obtained from $G$ by replacing $x$ with a clique. Then $(I(H))^{s}$ has linear quotients.
Remark 3.6.
If $G$ is a (gapfree) graph and $x,y$ are vertices such that $N_{G}(x)\setminus\{y\}=N_{G}(y)\setminus\{x\}$, then 2.3 or 3.4 can be used to simplify $G$ by identifying $x$ and $y$ prior to determining whether $I(G)^{s}$ has linear quotients for some $s$, potentially significantly reducing the number of generators that need to be ordered.
In the next two sections, we provide examples of gapfree graphs for which powers of their edge ideals have linear quotients starting from the second power. Applying the main results of this section to these graphs yields many additional examples of graphs whose powers have linear quotients.
@@ 4. Ordering of pure powers
Let $G$ be a graph with $E(G)=\{e_{1},\ldots,e_{s}\}$ and suppose that a power $I(G)^{q}$ of the edge ideal $I(G)$ has linear quotients (which particularly implies that $G$ is gapfree) with an ordering
${\mathcal{N}}^{(q)}:u_{1}>\cdots>u_{r}$
of the generators of $I(G)^{q}$, where pure powers appear in the ordering with
@@
${\mathcal{N}}^{(q+1)}:u_{1}e_{1}>\cdots>u_{r}e_{1}>u_{1}e_{2}>\cdots>u_{r}e_{2}>\cdots>u_{r}e_{s}$
of the generators of $I(G)^{q+1}$, where $u_{i}e_{j}$ is omitted if $u_{i}e_{j}=u_{k}e_{\ell}$ with $\ell<j$, is called the efficient ordering constructed from the ordering ${\mathcal{N}}^{(q)}$. Recursively, we define the efficient ordering ${\mathcal{N}}^{(s)}$, constructed from the ordering ${\mathcal{N}}^{(q)}$, of the generators of $I(G)^{s}$ for $s\geq q$.
The following is a refinement of Conjecture 1.2 by explicitly describing the potential linear quotient ordering for higher powers of $I(G)$ coming from that of a given power. It provides a strong means of approaching Conjecture 1.2
@@
Let $G$ be a graph and suppose that $I(G)^{q}$, for some $q\in{\mathbb{N}}$, has linear quotients with an ordering ${\mathcal{N}}^{(q)}:u_{1}>\dots>u_{r}$ of its generators. Then, $I(G)^{s}$ has linear quotients with the efficient ordering ${\mathcal{N}}^{(s)}$, constructed from ${\mathcal{N}}^{(q)}$, of its generators for all $s\geq q$.
@@
Let ${\mathcal{A}}(e_{1})$ denote the set of $w_{i}$ which can be written as a product of $q$ edges, one of which is $e_{1}$. In this case, we say $w_{i}$ has a presentation using $e_{1}$. Let ${\mathcal{A}}(e_{2})$ denote the set of $w_{i}\not\in{\mathcal{A}}(e_{1})$ which have a presentation using $e_{2}$, ${\mathcal{A}}(e_{3})$ the set of $w_{i}\not\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$ which have a presentation using $e_{3}$ and ${\mathcal{A}}(e_{5})$ the set of $w_{i}\not\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$ which have a presentation using $e_{5}$. Finally, ${\mathcal{A}}(e_{4})=\{e_{4}^{q}\}$.
We claim that for any $2\leq p\leq\binom{5+q-1}{q}$, the colon ideal $(w_{1},\ldots,w_{p-1}):w_{p}$ is generated by a subset of $\{a,b,c,d,e\}$. If $w_{p}\in{\mathcal{A}}(e_{1})$, then the claim follows from the induction hypothesis.
CASE 1: Let $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{2})$ with $i<p$. Let $w_{i}=u_{\ell}e_{2}$ and $w_{p}=u_{t}e_{2}$ with $s<t$. Then there is $u_{\ell}$ with $\ell<t$ for which $u_{\ell}:u_{t}$ is a variable and $(u_{\ell}:u_{t})|(u_{\ell}:u_{t})$. Set $w_{j}=u_{\ell}e_{2}$ and notice that $j<p$. Note that $w_{j}$ may belong to ${\mathcal{A}}(e_{1})$. Then $w_{j}:w_{p}$ is a variable and $(w_{j}:w_{p})|(w_{i}:w_{p})$, as desired.
CASE 2: Let $w_{p}\in{\mathcal{A}}(e_{2})$ and $w_{i}\in{\mathcal{A}}(e_{1})$. If $a$ does not divide $w_{p}$, then $w_{i}:w_{p}$ is divisible by $a=(e_{1}w_{p}/e_{2}):w_{p}$. If $e_{5}=ae$ appears in the presentation of $w_{p}$, then both $(e_{1}w_{p}/e_{5}):w_{p}=b$ and $(e_{1}e_{3}w_{p}/e_{2}e_{5}):w_{p}=d$ belong to $(w_{1},\ldots,w_{p-1}):w_{p}$. Let $w_{i}=e_{1}^{f_{1}}e_{2}^{f_{2}}e_{3}^{f_{3}}e_{4}^{f_{4}}e_{5}^{f_{5}}$ and $w_{p}=e_{2}^{f^{\prime}_{2}}e_{3}^{f^{\prime}_{3}}e_{4}^{f^{\prime}_{4}}e_{5}^{f^{\prime}_{5}}$. If one of the inequalities
@@
which is impossible, since $\sum_{i=1}^{5}f_{i}=\sum_{i=2}^{5}f^{\prime}$ and $f_{1}>0$.
CASE 3: Let $w_{p}\in{\mathcal{A}}(e_{3})$. Since $e_{2}(w_{p}/e_{3})\in{\mathcal{A}}(e_{2})$, it follows that $(e_{2}w_{p}/e_{3}):w_{p}=b$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $b$ does not divide $e_{p}$, if $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$, then $b$ divides $w_{i}:w_{p}$. If $w_{i}\in{\mathcal{A}}(e_{3})$, then a similar argument as in Case 1 is valid without modification.
CASE 4: Let $w_{p}\in{\mathcal{A}}(e_{5})$. Since $e_{1}(w_{p}/e_{5})\in{\mathcal{A}}(e_{1})$, it follows that $(e_{1}w_{p}/e_{5}):w_{p}=b$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $b$ does not divide $e_{p}$, if $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$, then $b$ divides $w_{i}:w_{p}$. If $e_{4}$ appears in the presentation of $w_{p}$, then $e_{3}(w_{p}/e_{4})\in{\mathcal{A}}(e_{3})$ and $(e_{3}w_{p}/e_{4}):w_{p}=c$ belongs to $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $c$ does not divide $e_{p}$, if $e_{i}\in{\mathcal{A}}(e_{3})$, then $c$ divides $w_{i}:w_{p}$. Let $w_{p}=e_{5}^{q}$. Since $e_{4}e_{5}^{q-1}>e_{5}^{q}$ and $e_{4}e_{5}^{q-1}:e_{5}^{q}=d$, then $d$ belongs $(w_{1},\ldots,w_{p-1}):w_{p}$. Since $d$ does not divide $w_{p}=e_{5}^{q}$, if $e_{i}\in{\mathcal{A}}(e_{3})$, then $d$ divides $w_{i}:w_{p}$. If $e_{i}\in{\mathcal{A}}(e_{5})$, then a similar argument as in Case 1 is valid without modification.
CASE 5: Finally, if $w_{p}\in{\mathcal{A}}(e_{4})$, in other words, $e_{p}=e_{4}^{q}$, then a routine computation done as above yields $(w_{1},\ldots,w_{p-1}):w_{p}=(a,c)$. This completes the proof. ∎
@@
5. Pentagon together with one vertex
Developing the argument from the proof of Proposition 4.2, we now study finite graphs formed by carefully adding vertices to the pentagon. We first study the case where a single vertex is added with specific connecting edges.